OpenStudy (anonymous):

evaluate the integral from negative infinity to infinity of (1 / (e^x + e^-x)) dx

OpenStudy (valpey):

\[\int_{x=-\infty}^{\infty}\frac{1}{e^{x}+e^{-x}}dx\]

OpenStudy (anonymous):

yes

OpenStudy (primeralph):